OpenStudy (anonymous):

(x^2y)^2 can you help me simplify it

Parth (parthkohli):

Is the question 1 or 2?\[1) \quad {(x^{2y})^2}\]or\[2) \quad (x^2 y)^2\]

OpenStudy (anonymous):

it number 2

Parth (parthkohli):

OK, both the questions use only one formula, given below:\[\left(uv \right)^2 \implies u^2v^2\]

OpenStudy (anonymous):

so both the exponents are 2

Parth (parthkohli):

\[\left(x^2 y \right)^2 \implies \left( x^2\right)^2 y^2\]

OpenStudy (anonymous):

so is the answer x^2 y^2

Parth (parthkohli):

Nope...it's \((x^2)^2y^2\) You can simplify \((x^2)^2\)

OpenStudy (anonymous):

but thats not one of my choices